OpenStudy (thcsbabygirl1224):

If you are riding on a Ferris wheel that is 60 feet tall and it takes you 5 minutes to make one revolution, what is your linear velocity? 0.62 feet per second 0.31 feet per second 37.68 feet per second 3.14 feet per second 9.42 feet per second i think it is d but not sure

OpenStudy (thcsbabygirl1224):

@AloneS @mathstudent55

OpenStudy (alones.):

There's two ways I can explain this. 1. d = 60ft C = (pi)d = 188.5ft 188.5ft/5min = 37.7ft/min 37.7ft/min/60sec/min = 0.63ft/sec (A)

OpenStudy (alones.):

2. First you find the circumference of the Ferris wheel, which is pi * diameter, which equals to 188.46 ft. So, now you know that the rate of speed is 188.46 ft. / 5min, but we need the rate of speed in feet per second, not feet per 5 minutes. So, we divide both sides by 5, 188.46 /5 = 37.692 and 5 / 5 = 1. Now, we have the rate of speed equal to 37.692 ft. / minute, but this is still not feet per second. Because there is 60 second in one minute, we divide both sides by 60. 37.692 / 60 = 0.62 and 1min / 60 = 1 second. So now we have the rate of speed equal to 0.62 feet / second.

OpenStudy (alones.):

So, the answer is... what?

OpenStudy (thcsbabygirl1224):

.62?

OpenStudy (alones.):

Yep!

OpenStudy (alones.):

A.

OpenStudy (thcsbabygirl1224):

thank you i will message you and ask questions im trying to try to pass

OpenStudy (alones.):

Okay, and no problem!!